Consider the following income statement for the Heir Jordan Corporation:

A 20 percent growth rate in sales is projected.
Prepare a pro forma income statement assuming costs vary with sales and the dividend payout ratio is constant. (Input all answers as positive values. Do not round intermediate calculations.)
  
What is the projected addition to retained earnings? (Do not round intermediate calculations.)
Consider the following income statement for the Heir Jordan Corporation: HEIR JORDAN CORPORATION Income Statement Sales Costs $49,000 40,300 Taxable income Taxes (22%) 8,700 1,914 Net income 6,786 Dividends Addition to retained earnings $2,400 4,386 A 20 percent growth rate in sales is projected. Prepare a pro forma income statement assuming costs vary with sales and the dividend payout ratio is constant. (Input all answers as positive values. Do not round intermediate calculations.) HEIR JORDAN CORPORATION Pro Forma Income Statement Sales Costs Taxable income Taxes Net income What is the projected addition to retained earnings? (Do not round intermediate calculations.) Addition to retained earnings
Answer #1

Answer::

Sales =(1+20%)*49000=$58800

Cost=(1+20%)*40300=$48360

Income Statement
Sales  S 58800
Costs  C 48360
Taxable income EBT=S-C 10440
Taxes T=22%*EBT 2296.8
Net Income N=EBT-T 8143.2

Earlier Dividend Ratio =2400/6786=35.37%

So Additional retained earning =(1-35.37%)*8143.1=$5263.20